OpenStudy (anonymous):

a line passes through the point (6,8) and has a slope of (2). write an equation for this line

OpenStudy (anonymous):

y = 2x + b 8 = 2(6) + b 8 = 12 + b b = -4 y = 2x - 4

OpenStudy (anonymous):

danm your fast moneybird. thank you

OpenStudy (anonymous):

Glad that you get help

OpenStudy (anonymous):

@moneybird, i need help with a graph. is there a way i can email it to you? Please